In a survey of 100 randomly selected professional soccer players, the mean counter movement jump height was x=55.08cm. Suppose that somehow we knew that the population standard deviation of the counter movement jump is 5.52 cm.

Find a 90% confidence interval estimate of the mean CMJ of all professional soccer players.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 55.08 cm

Population standard deviation = \sigma = 5.52 cm

Sample size = n = 100

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.1 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (5.52 / \sqrt 100)

= 0.91

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

55.08 - 0.91 < \mu < 55.08 + 0.91

54.17 < \mu < 55.99

(54.17 , 55.99)